### v1

# K-Edit: Language Model Editing with Contextual Knowledge Awareness

Abstract
As the world changes, we need to be able to update our models and correct false information without costly retraining. Knowledge-based model editing enables precise modifications to the weights of large language models in order to modify the information encoded within. Recent approaches have seen success in enabling recall of edited information for thousands of edits at once. However, these approaches fail to produce edits that account for associated contextual information. We present K-Edit, an effective approach to generating contextually consistent knowledge edits. By using knowledge graphs, which maintain contextual consistency when an edge is edited, we are able to generate additional contextual edits that ensure consistency of related information in the language model. Our experiments demonstrate significant improvements in multi-hop question answering while maintaining the general effectiveness and scalability of model edits.
1 Introduction
Figure 1: Example of how K-Edit improves model editing. Whereas existing methods edit the model so that it recalls the correct updated information, that new information is not associated with contextual information. The result is that the LLM is left with an inconsistent world model in which “The Prime Minister is Rishi Sunak”, and “Rishi Sunak is married to Akshata Murty”, but “The Prime Minister’s spouse is Carrie Johnson”. K-Edit fixes this by querying a knowledge graph (such as Wikidata) for contextual information, and then turning that information into contextual edits, and applying it to the LLM to improve multi-hop reasoning and consistency.
The ability to pinpoint and edit knowledge within the parameters of a Large Language Model (LLM) is a powerful tool for maintaining alignment with an ever-changing world. Existing approaches to direct model editing make shallow edits that fail to integrate with contextual knowledge the model possesses (Zhong et al. 2023). This work proposes and explores a simple solution to better enhance knowledge reasoning over edits by including contextual and associated knowledge in the editing process. We use the contextual consistency of knowledge graphs when edited to improve the consistency of language models under direct model editing.
Direct Model Editing (or Knowledge-based Model Editing) aims to change the implicit knowledge in an LLM’s weights with no other changes to how the model is used (Wang et al. 2023b; Mazzia et al. 2023; Yao et al. 2023). Prior methods have focused on the ability to edit and recall a new fact without affecting neighboring facts or general generation ability (Zhu et al. 2020; Mitchell et al. 2021; Meng et al. 2022a, b). Recently, however, researchers have begun to question the effectiveness of these edits through benchmarks evaluating the ability of models to use these edits for multi-hop reasoning (Zhong et al. 2023). For example, if we edit the British Prime Minister to be Rishi Sunak instead of Boris Johnson, we would expect the model to be able to answer Who is the British Prime Minister’s wife? with the correct answer Akshata Murty. Existing direct model editing techniques largely fail in this regard. We address this issue by utilizing the relational information in knowledge graphs.
In this work, we propose K-Edit, a simple, yet effective, approach of adding more contextual knowledge into the editing process. In contrast to text (and thus LLMs), knowledge graphs are very simple to edit. Knowledge graphs are graph databases encoding the relations between real-world entities. Since the structure of knowledge is explicit in a knowledge graph, editing a single edge immediately updates connected associations as well. We propose a method that incorporates these automatic associations into the editing process. After an edit is made, we extract related contextual information from a knowledge graph. We then construct contextual edits that ensure the context is ingrained in addition to the initial edit. Figure 1 shows the advantage of this approach.
Our results show that this method significantly helps with the primary goal, as measured by significant improvement on multi-hop reasoning over edited information. We even show that such edits improve multi-hop reasoning even when the required reasoning path does not directly appear in the context. Our results also indicate that there is no cost to quality as measured by other knowledge editing metrics, and that thousands of such edits can be applied without degradation compared to baselines.
Our contributions are
  1. 1.
    K-Edit, a new direct model editing algorithm for improving contextual and associated knowledge awareness through incorporation of knowledge graph data.
  2. 2.
    Experimental evaluation that shows improved multi-hop reasoning with K-Edit and its ability to generalize.
  3. 3.
    Additional experiments and ablations showing no degradation to other knowledge editing metrics and the ability to apply thousands of such edits at a time.
2 Background
2.1 Knowledge Graphs
[…]
2.2 Knowledge-based Model Editing
We define some background notation and concepts for knowledge-based model editing. Here we deal with edits of the form, $(s,r,o\rightarrow o^{*})$, meaning we want to edit subject $s$ such that the relationship $r$ now refers to object $o*$ instead of $o$. We define a text prompt $t_{r}(s)$ for the edit (also referred to as a cloze statement) in which $o^{*}$ is the desired output, e.g. $t_{\text{prime minister}}(\text{{United Kingdom}})$ = “The Prime Minister of {the United Kingdom} is”. In this example the United Kingdom is the entity for which we want to modify the Prime Minister association. In general, knowledge-based model editing approaches seek to modify the model parameters $\theta$ such that the prompt $t_{r}(s)$ will lead to the generation of text $o^{*}$.
2.3 Batch Editing
In order for our approach, K-Edit, to succeed, we rely on the ability to make edits in a batch. Our approach works with any such algorithm. For our experiments, we use MEMIT, a state-of-the-art batch editing algorithm that can scale to thousands of edits at once (Meng et al. 2022b). MEMIT is the best option to use with K-Edit, and the only option that can scale to large numbers of edits.
MEMIT has three key properties: (1) it computes the desired model parameter changes for thousands of edits independently, (2) it precisely targets associations of the subjects to be edited, and (3) it can apply all edits together without significant model degradation. Further background details on how MEMIT works can be found in Appendix C as well as in their paper.
3 K-Edit Algorithm
K-Edit leverages knowledge graphs to improve the quality of model edits by incorporating contextual information. The algorithm consists of three main steps that work together to create more comprehensive knowledge updates.
Existing approaches modify representations of the subject $s$ such that the model generates the desired association $o^{*}$ for relationship $r$. However, generating the associated text is not the same as internalizing the association. K-Edit uses external knowledge bases to improve the quality of model edits such that the new subject representation is integrated with connected contextual associations.
K-Edit works according to the following steps. (1) initial edits are applied the model using a batch editing approach (e.g. MEMIT). (2) Contextual knowledge is extracted from a knowledge graph around the edit and converted to contextual edits. (3) The contextual edits are applied using the same batch editing approach from step 1. This can be repeated for deeper depths of multi-hop contextual knowledge. Algorithm 1 shows the high level algorithm for implementing K-Edit.
Algorithm 1 K-Edit Algorithm
0:  Model $\theta_{0}$, Batch of initial edits $B$, KG $\mathcal{K}=(\mathcal{E},\mathcal{R},\mathcal{T})$, context depth $k$
1:  $\theta_{1}\leftarrow\text{BatchEdit}(\theta_{0},B)$
2:  for $i\in[2,...,k]$ do
3:     $C\leftarrow\text{GetContextualEdits}(B,\mathcal{K},i)$ edits of depth i
4:     $\theta_{i}\leftarrow\text{BatchEdit}(\theta_{i-1},C)$
5:  end for
6:  return  $\theta_{k}$ Returns the final model parameters
MQuAKE MultiCounterfact
#examples 3,000 10,000
#relations 36 34
Table 1: Dataset statistics.
1-edit 2-edits 3-edits 4-edits
2-hops 513 487 NA NA
3-hops 356 334 310 NA
4-hops 224 246 262 268
Table 2: Breakdown by number of hops and number of edits for the MQuAKE dataset (MQuAKE-CF-3k version)
3.1 Creating Contextual Edits
Algorithm 2 GetContextualEdits
0:  Batch of initial edits $B$, KG $\mathcal{K}=(\mathcal{E},\mathcal{R},\mathcal{T})$,current context depth $d$
1:  $C_{1}\leftarrow B$
2:  for $i$ in $[2,...,d]$ do
3:     $C_{i}\leftarrow\emptyset$ Contextual edits of depth i
4:     for all $(s,r_{1},...,r_{i-1}\rightarrow o_{i-1}^{*})\in C_{i-1}$ do
5:        # Next-hops based on previous depth’s target object
6:        $N\leftarrow\{(s_{i},r_{i},o_{i})\in\mathcal{T}\,|\,s_{i}=o_{i-1}^{*}\}$
7:        for all $(o_{i-1}^{*},r_{i},o_{i}^{*})\in N$ do
8:           # Construct new edits using the next-hop edges
9:           $e\leftarrow(s,r_{1},...,r_{i-1},r_{i}\rightarrow o_{i}^{*})$
10:           $C_{i}\leftarrow C_{i}\cup\{e\}$
11:        end for
12:     end for
13:  end for
14:  $O\leftarrow\emptyset$
15:  for all $(s,r_{1},...,r_{d}\rightarrow o_{d}^{*})\in C_{d}$ do
16:     # Create text prompts for each edit of depth d
17:     $O\leftarrow O\cup\{t_{r_{d}}(r_{d-1}(...(r_{1}(s)))\rightarrow o_{2}\}$
18:  end for
19:  return  $O$# Returns all contextual edits of depth d
Contextual edits are model editing prompts that are constructed to not just reinforce a simple edited fact, but also to ensure the fact is linked with connected facts (Examples of connected facts in figure 1). With this goal in mind, we design contextual edits that both require the model to recall the edited information on its own and require the model to follow the multi-hop relationships. The following section describes the construction of such prompts.
[…]
We also note that we treat the United Kingdom Prime Minister as the subject to edit which means the [Minister] token rather than [Kingdom] token is the one that is the target of the update. We do this in accordance with the findings of Geva et al. (2023) that the final token of a subject aggregates the information and relationships from preceding tokens. This process allows the model to learn the 2nd hop association in the context of the previous edit as opposed to memorizing all 2-hop associations for the subject token. We note that this setup is coincidentally similar to the approach used to analyze models in Yang et al. (2024). It is encouraging that our approach to edit multi-hop associations corresponds to the way they discovered models perform latent multi-hop reasoning in the first place. Algorithm 2 presents the detailed description for constructing contextual edits of any hop depth.
3.2 Scaling Contextual Edits
One of the challenges with generating multi-hop contextual edits is that the number of templates for editing rapidly explodes. For instance, with 50 relationship types, there are potentially 2,500 different templates required for 2-hop contextual edits (though most of those will not exist in practice). This could be solved naively by generating the templates with an LLM but that could still get burdensome for larger KGs or hop depths and could introduce errors. Instead, our approach employs a linguistic mechanism that enables the scalable creation of contextual edits for any hop length.
To use this method all that is needed is two text templates for each relation type $r$. We require one template $t_{r}(s)$ that creates a cloze statement for each relation $r$. This is a fill-in-the-blank style prompt. For example, the template “The Prime Minister of {s} is” prompts the model to complete the association. Additionally we need one more template, $r(s)$ that creates a subject-relation clause that semantically refers to the corresponding object e.g. ”the spouse of {s}”. With these mappings we create a single cloze statement template for a multi-hop edit, using the template $t_{r_{1}}(r_{2}(...(r_{k}(s)))$. This means we only need to create $2\times|\mathcal{R}|$ templates rather than $|\mathcal{R}|^{k}$. Note that their is no dependence on depth $k$. Algorithm 2 presents the method for constructing these contextual edits. These more simple templates enable the easy use of LLMs for further generation (Appendix E).
4 Experiments
We run experiments on K-Edit to evaluate its ability to incorporate contextual information under edits as well as the general quality of the model edits produced.
4.1 Datasets
We use two datasets, MQuAKE (Zhong et al. 2023) and MultiCounterFact(Meng et al. 2022b). Dataset statistics can be found in Table 2.
The primary dataset, MQuAKE, is used for evaluating multi-hop reasoning on edits (Zhong et al. 2023). This dataset contains 3,000 multi-hop questions (2-4 hops) with each answer based on 1-4 counterfactual edits. For example, the question “What is the headquarters location for the broadcaster of True Detective?” is paired with two edits: (True Detective, broadcaster, HBO $\rightarrow$ ABC) and (ABC, headquarters location, New York City $\rightarrow$ Los Angeles). These edits are not real in the world, but are plausible counterfactual edits that could be applied. This example was a two-hop question with two edits, however, the number of edits does not need to be the same as the number of hops. The exact breakdown can be found in Table 2.
We also use a subset of MultiCounterFact dataset (Meng et al. 2022b) derived from CounterFact (Meng et al. 2022a). This dataset contains 30,000 counterfactual edits and is used to evaluate the strength of an edit, whether unaffected neighboring relations remain unchanged, and the generation ability of the model post editing. The MultiCounterFact modification ensures there are no direct conflicts between edits in the dataset (i.e. no duplicate edits for a given subject and relation). We use a subset of 10,000 questions for our experiments.
Model Algorithm Multi-Hop Multi-Hop CoT
[…]
K-Edit 14.5 (0.64) 17.0 (0.63)
Falcon-7b Base 38.1 43.3
MEMIT 5.87 12.57
K-Edit 11.1 (0.57) 18.43 (0.71)
Table 3: Results on MQuAKE multi-hop question answering and reasoning dataset over edited information. Results show that K-Edit significantly improves Multi-Hop QA and is similar to ROME on chain-of-thought question answering. Base gives the results of the unedited model on the original answers. Parenthesis show standard error of the mean estimate. $\dagger$ indicates results from Meng et al. (2022b)
Model Algorithm Multi-Hop Multi-Hop CoT
[…]
GPT-J-6B K-Edit (Generalize) 11.5 (0.58) 14.8 (0.65)
K-Edit 14.5 (0.64) 17.0 (0.63)
MEMIT 5.87 12.57
Falcon-7b K-Edit (Generalize) 8.2 (0.50) 16.3 (0.68)
K-Edit 11.1 (0.57) 18.4 (0.71)
Table 4: Generalization experiment in which the multi-hop edge being evaluated is removed from the contextual edits. This shows that some of the K-Edit improvement stems from improved representations rather than more broad memorization of associations. Results are for MQuAKE dataset. Parenthesis show standard error of the mean estimate. $\dagger$ indicates results from Meng et al. (2022b)
4.2 Baselines
We compare against a number of alternative direct model editing baselines.
FT-W is a finetuning approach with weight decay to avoid altering other memories or changing unrelated model behavior (Zhu et al. 2020).
MEND uses a hypernetwork meta-trained to modify fine-tuning gradients so that they produce higher quality edits (Mitchell et al. 2021).
ROME uses counterfactual tracing to locate what layer of the model contains the information to be edited before updating the MLP component in that layer (Meng et al. 2022a).
MEMIT (presented above) uses a similar objective to ROME but modified to enable editing large sets of facts at once, which none of the other baselines are able to do well (Meng et al. 2022b). We consider this to be the main baseline as it is also the editing mechanism we use in our K-Edit implementation.
While there are a number of knowledge editing approaches based on retrieval mechanisms, these methods fundamentally alter the nature of the model. Our focus is on direct model editing approaches. This is discussed further in the related works section.
5 Results and Analysis
Our results show that K-Edit improves multi-hop contextual reasoning over edited information without sacrificing in terms of scalability or individual edit performance.
Table 3 shows our main results for GPT-J and Falcon-7b on the MQuAKE dataset. We see that the model’s multi-hop answering ability vastly exceeds the other baselines with a relative improvement of 57.6% over the best baseline MEND (14.5% versus 9.2%) and close to doubling the performance of ROME and MEMIT (+90.8% and +79.0% respectively). Under Chain-of-Thought prompting, K-Edit significantly outperforms MEMIT while slightly underperforming ROME. We note that the MQuAKE results are from applying the edits for one question at a time. ROME has noted weaknesses when applying edits in bulk as will be discussed later.
We also compare K-Edit to MEMIT for Falcon-7b and find a even greater improvement than for GPT-J with the multi-hop accuracy nearly doubling.
While K-Edit significantly improves multi-hop accuracy, we note that there is still room for improvement. The GPT-J Base model gets 43.4% accuracy on the unedited answers which sets a rough upper bound for where we could expect to see results.
Algorithm Num Hops
2 3 4
[…]
K-Edit (Generalize) +1.9 +5.2 +4.2
K-Edit (Memorize) +6.0 +0.3 +0.4
Table 5: Breakdown by number of hops showing the incremental accuracy gain over MEMIT due to generalization and memorization. The gain from generalization is spread across all the question types. While as expected the gain from memorization is concentrated on the two hop questions. Results are for GPT-J-6B on the MQuAKE dataset using CoT reasoning.
Algorithm Score Efficacy Paraphrase Specificity Fluency Consistency
FT-W@1 64.4 100.0 73.8 43.4 614.7 34.9
MEND@1 62.8 97.2 52.8 53.9 620.6 32.3
ROME@1 91.9 99.9 99.5 79.5 619.9 42.5
MEMIT@1 91.6 99.8 95.2 81.7 621.5 41.5
K-Edit@1 90.1 100.0 (0.0) 91.6 (0.85) 80.8 (0.98) 622.2 (0.53) 40.7 (0.44)
FT-W@1,000 71.1 99.6 75.2 53.1 258.1 17.0
MEND@1,000 48.4 45.2 46.3 54.9 527.4 15.4
ROME@1,000 60.3 68.2 64.4 51.0 509.9 6.8
MEMIT@1,000 90.5 99.7 93.4 80.6 622.0 41.4
K-Edit@1,000 90.6 99.8 (0.14) 95.3 (0.56) 79.5 (0.81) 620.2 (0.55) 41.7 (0.39)
Table 6: Results on MultiCounterfact dataset (from MEMIT) with GPT-J. K-Edit does not degrade and in fact, may actually improve editing ability. Results@1 indicates performing a single edit at a time. @1,000 indicates performing 1,000 edits at once. Efficacy measures memorization of the edit. Paraphrase measures robustness to paraphrases. Specificity measures whether other similar facts remain unchanged. Fluency measures entropy of generated text. Consistency measures the TF-IDF similarity between generated text about the subject and a reference text about the new object. Formal definitions are in the appendix.
Algorithm Multi-Hop Multi-Hop CoT
MEMIT 8.1 12.3
K-Edit-Parallel 12.6 (0.61) 12.7(0.61)
K-Edit 14.5 (0.64) 17.0 (0.63)
Table 7: Results showing the necessity of sequential updates in K-Edit.
Multi-Hop Accuracy
[…]
1 30.6 7.0 28.1 22.4
2 15.6 9.3 16.67 13.9
3 NA 2.3 5.0 3.5
4 NA NA 8.2 8.2
All 23.3 6.3 13.9
Multi-Hop CoT Accuracy
[…]
1 31.0 5.6 20.5 20.6
2 20.9 7.8 21.5 17.0
3 NA 18.4 12.2 15.6
4 NA NA 6.0 6.0
5 26.1 10.3 14.7
Table 8: GPT-J multi-Hop accuracy on MQuAKE by number of hops and number of edits.
5.1 Generalization
One question that may arise is whether the benefit of K-Edit is simply from the multi-hop questions being covered by the contextual edits. Specifically, are the observed gains due to brute force inclusion of related information or is there an aspect of K-Edit that improves the edited representations in a more generalized way? To address this question, we run additional ablation experiments in which the question edges from MQuAKE are filtered out of the contextual edits. We present these experiments in Table 4. We see that while performance does decrease (as expected), it is still significantly better than the original MEMIT performance. Specifically, we observe that roughly half of the improvement from K-Edit generalizes to edges not covered by contextual edits. This is a somewhat surprising result. Including contextual edits on tangentially related edges improves the quality of original edit such that other multi-hop information is easier for the model to retrieve. The implication of this is that the quality of K-Edit is not solely limited to the relation types present in the knowledge graph.
When we break down these gains by the number of hops, we see that the gains from generalization are spread across the dataset, with gains on 2, 3, and 4-hop questions. Table 5 shows the incremental gain from generalization and from memorization on GPT-J using CoT reasoning. We see that the incremental gain from memorized associations are concentrated on the 2-hop questions. This is expected since we only use 2-hop contextual edits in our experiments. This is also evidence that K-Edit could perform even better if contextual edits of greater depth were included.
When we filter the analysis for questions with only a single edited hop, we find only a 2.8% accuracy difference performance between full K-Edit and the generalization ablation, and absolutely no difference if employing CoT reasoning. This makes sense since questions with multi-edits would be especially hard to generalize. Generalization relies on using some degree of the LLM’s implicit knowledge. When that knowledge is also edited, it would be much harder to associate the correct contextual information.
5.2 Edit Quality Remains High for Standard Metrics and Larger Scales
K-Edit’s edit quality remains high when evaluated under metrics other than multi-hop accuracy (see Table 6). These experiments show the results of running editing algorithms when only a single edit is applied at a time (@1) or when batches of 1,000 edits are applied at a time (@1,000). We see that K-Edit is at or near state-of-the-art for all metrics for both @1 and @1,000. While ROME slightly outperforms K-Edit in @1 results, it dramatically underperforms in results @1,000. The overall score (harmonic mean of efficacy, paraphrase, and specificity) for ROME@1,000 is 60.3 relative to K-Edit’s 90.6.
Compared to MEMIT we see highly comparable scores across the board. We note that while K-Edit does introduce additional edits in the form of contextual edits, these edits are primarily there to reinforce the associated information of the initial edit. As a result, there is no noticeable impact to scalability. Though further research is needed to understand the full scalability limits of K-Edit.
[…]
5.3 Sequential Ordering is Important
K-Edit applies the contextual edits after applying the original edits. This is so that the contextual edits can refine the initial edit to incorporate the related information. We find that running K-Edit in parallel (meaning the model updates for both original edits and contextual edits are computed at the same time) leads to significant decline in performance. Table 7 shows the importance of using sequential edits compared to parallel ones (14.5 to 12.6 and 17.0 to 12.7 for CoT reasoning). We theorize that when the contextual edits are not applied after the initial edits, the model may be updating the wrong associations. For instance, in our running example, if Boris Johnson is still associated as “the United Kingdom Prime Minister”, the algorithm may be updating the spouse of Boris Johnson rather than enhancing the association of Rishi Sunak as Prime Minister.
5.4 Analysis by number of hops and edits
We see an overall trend of declining performance with increasing number of hops. However, the 4-hop questions have actually better performance than the 3-hop questions. The primary reason for this is that the MQuAKE dataset contains a more limited variety of reasoning paths for 4-hops relative to 3-hops and the LLM happens to perform better with those style of questions.
[…]
6 Related Works
Direct Model Editing Works
These approaches include aforementioned works (Zhu et al. 2020; Mitchell et al. 2021; Meng et al. 2022b, a). Other works use a hypernetwork to predict parameter updates for a model Cao, Aziz, and Titov (2021)
Retrieval Based Works
Outside of direct model editing approaches (Zhu et al. 2020; Meng et al. 2022a; Mitchell et al. 2021; Meng et al. 2022b; Ma et al. 2023), other approaches have sought to adapt to changing knowledge via storing external edit memories and employing various in-context learning and retrieval mechanisms. IKE and others add all edits in the LLM context (Zheng et al. 2023; Onoe et al. 2023); SERAC identifies if a related fact has been edited and retrieves it (Mitchell et al. 2022); MeLLo and DeepEdit iteratively retrieve edited memories during the generation and decoding process (Zhong et al. 2023; Wang et al. 2024).
We note that these retrieval-based editing approaches are fundamentally different than direct model editing approaches. Direct model editing leaves the model architecture and code entirely the same with the only change being to parameter weights. This means any system developed to use the original model can use the updated model with no other changes. In contrast, retrieval-based editing systems fundamentally change how the model is used and functions. For example, MeLLo requires setting up a retrieval database, iteratively prompting the LLM, making multiple calls to the database, and prompting the LLM to create further sub-questions (Zhong et al. 2023). This fundamentally changes the characteristics of the model, increases latency, and alters model behavior for non-QA tasks. An alternate line of work aims to supplement the parametric knowledge of LLMs by augmenting the model prompt with information retrieved from an indexed corpus (Lewis et al. 2020) or Knowledge Graphs (Markowitz et al. 2024).
Additional Aspects of Model Editing
Other papers have looked into other aspects of contextual knowledge for edits such as reversibility (Ma et al. 2023), multi-linguality (Wang, Haddow, and Birch 2023; Si, Zhang, and Zhang 2024), or other implied logical conditions given an edit (Wang et al. 2023a; Cohen et al. 2023; Li et al. 2023).
7 Future Work
Future directions include both straightforward advancements as well as further analytical directions. Simple advancements include extending K-Edit to broader and deeper contexts. This can include more relation types, increasing the number of iterations to $k>2$, and including contextual edges connected to the edit subject $s$ instead of just $o^{*}$. These directions could yield simple improvements and benchmark gains, especially on the MQuAKE dataset.
Other interesting future directions include counterfactual knowledge tracing for contextual information as well as better measures of contextual knowledge in edited representations beyond output related metrics. Further, more study on enhancing the generalization effect would enable the creation of better knowledge edits. While K-Edit makes meaningful advancement in contextual integration for knowledge editing, there is still more progress to be made.
8 Conclusion
This paper presents a K-Edit, an algorithm for improving knowledge-based model editing to better incorporate contextual information. Our experiments show that our approach greatly improves results on multi-hop question answering tasks when edits have been applied. We further show intriguing results demonstrating the potential for better contextual generalization of edits under K-Edit. In addition, our approach scales to thousands of edits at once with no significant degradation on standard metrics. These combine to show that K-Edit is a strong and easy tool for updating and editing knowledge in LLMs while maintaining consistency of information.
9 Limitations
Our method and experiments have a number of limitations. Our experiments are limited to adding context from a single additional hop from the edit target $o^{*}$, and only in the ”forward” direction. However, the evaluation contains up to four hops and also uses context from edges connected to the subject $s$ as well. This means that K-Edit could show improved results if subject context and deeper context were used.
We are also limited by the available datasets. Currently, there are only English based datasets for evaluating knowledge editing in multi-hop settings. In addition, we do not know if the multi-hop templating would function the same way in all languages. However, this could be mitigated with falling back to LLM-based generations if needed.
Finally, this evaluation is limited to 7 billion parameter models or less. While model editing approaches generally become more effective on more capable models, they can also become more computationally costly.
10 Ethical Statement
We find no major negative ethical implications of this work. Hopefully, enabling high quality model editing work will allow providers to avoid retraining models from scratch, thus avoiding costs and reducing environmental impact.
As noted in the limitations, this work is limited to English. We hope that the construction of multi-lingual model editing and multi-hop model editing datasets will enable this work to be studied under other languages. Since many knowledge graphs, Wikidata included, are multi-lingual, we see no reason why this work could not be applied in other languages.
Appendix A Hardware
The hardware requirements depend more on the underlying editing software, in this case MEMIT. We run software using 4 V100-16GB GPUs and make some optimizations for using multiple-GPUs and improving memory management.
Appendix B Definitions for Additional Edit Metrics
These definitions come from prior work and are used in Table 6.
Efficacy is the expectation that the target generation is more likely than the original fact.
Paraphrase is the expectation that the target generation is more likely than the original fact under paraphrased prompting.
Specificity is the expectation that neighborhood prompts remain correct after an edit. Neighborhood prompts are prompts for semantically related subjects instead of $s$.
Fluency is the weighted sum of the entropy of bi- and tri-gram distributions of the generated text.
Consistency is computed by generating text about $s$ and checking the TF-IDF similarity with a reference Wikipedia text about $o^{*}$. Specifically measured as cosine similarity between the TF-IDF vectors.
Appendix C MEMIT
MEMIT (Meng et al. 2022b) works as follows. MEMIT determines a small change to the $L$-th layer activation of the last token of $s$ (e.g. [Kingdom] in the example above) such that $t_{r}(s)$ produces the desired output $o^{*}$. It then modifies a subset of the model weights to produce the desired activation for a whole batch of such edits.
Let $h^{L}$ be the $L$-th activation of the last token of $s$. The target change is $z^{L}=h^{L}+\delta$ where $\delta$ is optimized to according to the following:
$\delta=\operatorname*{\arg\!\max}_{\delta}\frac{1}{P}\sum_{j=1}^{P}\log\mathbf{P}_{\theta(h^{L}+=\delta)}[o^{*}|x_{j}\oplus t_{r}(s)]$ (1)
This calculates a residual vector to be added to $h^{L}$ such that when $t_{r}(s)$ is input, the model outputs $o^{*}$ with maximum probability. $\theta(h^{L}+=\delta)$ represents the LLM in which $\delta$ is added to $h^{L}$. A weight decay penalty is applied to $\delta$ to minimize the change. $x_{j}$ are various contextual inputs pre-pended to $t_{r}(s)$ to improve generalization.
After computing each desired $\delta_{i}$ for each memory $i$ to be batch edited, MEMIT then computes how to spread those changes across a set of ”critical layers” $[l_{0}:L$]. Each $\delta_{i}$ is split into a residual $r_{i}$ to be added to each layer according to $r_{i}=\frac{\delta_{i}}{L-l+1}$. This assigns each of the critical layers an approximately equal amount of modification towards the desired $\delta_{i}$. The following optimization is then applied to the output matrix of the Multi-Layer Perceptron (MLP) block for each critical layer:
$\begin{split}\hat{W}_{l}^{out}=\operatorname*{\arg\!\min}_{\hat{W}}&\Biggl{(}\sum_{i=1}^{n}\left\|\hat{W}k_{i}-W_{l}^{out}k_{i}-r_{i}\right\|^{2}\\ &+\sum_{i=n+1}^{n+u}\left\|\hat{W}k_{i}-W_{l}^{out}k_{i}\right\|^{2}\Biggr{)}\end{split}$ (2)
Where $\hat{W}_{l}^{out}$ is the new final weight matrix of the MLP block of layer $l$, $W_{l}^{out}$ is the original weight matrix, $k_{i}$ is the activation input to the weight matrix being changed (intermediate activation of the MLP block). This computation modifies $W_{l}^{out}$ so that $n$ edited associations are modified by residual $r_{i}$ while not affecting $u$ unchanged associations. The actual computation is done using a more efficient format that does not require recomputing large numbers of unchanged associations. Further details can be found in the original paper.
Appendix D Hyperparameters
We introduce no additional hyperparameters beyond hop-count and use hyperparameters from Meng et al. (2022b) otherwise.
Appendix E Creating New Relation Templates with LLMs
The composable structure of our templates for contextual edits allows us to scale K-Edit to exponentially many combinations of multi-hop relations from the number of base templates we have. We note that this approach easily meets the needs of our experiments.
In actual deployments it may be worthwhile to include methods for scaling the number of relation types even further. Many knowledge graphs include thousands of different relation types. We introduce an LLM based approach to automatically constructing these base templates. The prompts we use can be found in figures 2 and 3.
Figure 2: The above prompt can be used to create relation templates of the form $t_{r}(s)$. These are cloze statements in which the next tokens would match the entity referred to from the subject entity {subject} using relation {relation}. To use this prompt, “{relation}” would be replaced with relation $r$.
Figure 3: The above prompt can be used to create relation templates of the form $r(s)$. These are noun-phrase type statements in which the referenced entity object has relation {relation} from subject {subject}. To use this prompt, “{relation}” would be replaced with the desired relation $r$.